find two consecutive negative integers whose product is 48 more than three times the first integer
Answer by CubeyThePenguin(3113)

consecutive negative integers: x, (x+1)

x(x+1) = 48 + 3x
x^2 + x = 48 = 3x
x^2 - 2x - 48 = 0
(x - 8)(x + 6) = 0
x = 8, x = -6

x is negative ---> integers: -6, -5